# TRG-Net: An Interpretable and Controllable Rain Generator

### v1

@@ 5.1 Rain Generator Verification
5.2 Unpaired Rain Generation
@@
Table 1: The FID and KID of unpaired rain generation on four rain datasets. Bold indicates the best result.
@@
Figure 8: The results of unpaired rain generation by VRG-Net and the proposed TRG-Net on three synthetic datasets and a real dataset. (a)-(d) are the rainless background images, the rainy images generated by VRG-Net, the rainy images generated by the proposed TRG-Net and the reference rainy images which are used for unpaired training, respectively. Every dataset displays two generated samples.
Table 2: The quantitative results of all competing methods on synthetic and real datasets. $*{}^{*}$ indicates the deraining results of PReNet [32] trained on the pseudo-paired data generated by method A. The best result is highlighted with bold.
@@
CycleGAN [$*{}^{*}$ 25.11 0.832 22.62 0.908 26.88 0.937 22.71 0.866
DerainCycleGAN [$*{}^{*}$ 28.50 0.910 21.10 0.871 22.31 0.938 22.01 0.863
DCD-GAN [$*{}^{*}$ 22.48 0.821 19.84 0.850 23.55 0.903 19.81 0.786
VRG-Net [$*{}^{*}$ 27.84 0.855 21.94 0.768 21.08 0.906 34.81 0.931
$*{}^{*}$ 32.33 0.949 24.86 0.959 32.15 0.983 35.80 0.945
Table 3: The segmentation performance of all competing methods on Cityscapes. Bold is the best result.
Methods mIoU mAcc
Rain 37.30 50.61
DSC [49] 28.82 39.12
JCAS [50] 43.74 57.35
CycleGAN [16] 40.85 54.13
DerainCycleGAN [21] 48.61 61.36
DCD-GAN [22] 29.89 40.57
CycleGAN [$*{}^{*}$ 50.86 64.38
DerainCycleGAN [$*{}^{*}$ 48.74 64.96
DCD-GAN [$*{}^{*}$ 44.21 60.02
VRG-Net [$*{}^{*}$ 36.99 50.28
$*{}^{*}$ 66.57 76.85
Clean (Upper bound) 72.50 80.82
@@
Table 4: The deraining results on synthetic datasets. Baseline means the derainers trained on the original dataset without augmentation. VRG-Net and TRG-Net denote augmented training using VRG-Net and the proposed TRG-Net, respectively. Bold is the best result.
@@
5.2.1 Unpaired Rain Generation
Experiment settings. We perform unpaired rain generation on three synthetic rain datasets, Rain100L, Cityscapes [30] and Kitti777 The rains in Cityscapes and Kitti are synthesised [30] with varying rainfall rate to evaluate the effect of rain on computer vision tasks in the outdoor environment, which consider the imaging mechanism, such as fog-like attenuation and rain photometry, under the condition of camera motion. For them, we select the 100mm/hr rain as training set and test set in this experiment. As the rain orientations in Cityscapes and Kitti are correlated with the position, we have further included three learnable masks on the rain maps to distinguish different orientations in different positions for the two datasets. [30], and a real rain dataset, SPA-Data [25], to verify the superiority of our rain generator. Specifically, to construct a complete unpaired training dataset, we remove the rainless images of half of the data samples and remove the rainy images of the other half of the data samples. The rotatable TV regularizer is used for the training of Rain100L. The generator is trained for 400 epochs on Cityscapes and Kitti, and 30 epochs and 250 epochs on Rain100L and SPA-Data, respectively. The same patch-based discriminator [17] is employed for all generators. Inspired by [52], we use the high frequency components of the rainy images instead of the rainy image itself as input to the discriminator for better discrimination. The other training settings are the same as those used in Sec. 5.1. We utilize Fréchet Inception Distance (FID) and Kernel Inception Distance (KID) as metrics for assessing the quality of generated rainy images, where lower values indicate better performance.
Experimental results. Table 1 shows the FID and KID results on four datasets. Our model achieves lower FID and KID compared to VRG-Net on all datasets, expecting the KID on Kitti, which means that the rainy images generated by the proposed TRG-Net has a higher rain quality than VRG-Net [18] and is able to capture the rain generation mechanism under different environment conditions. In Fig. 8, we show some typical results of the unpaired rain generation. It is easy to observe that the rain streaks generated by VRG-Net, which represents the CNN-based method, could be largely lacking in diversity, for example, the rain streaks in two random generation results on different background images are very similar, as shown in Fig. 8 (b). In contrast, the rainy images generated by the proposed TRG-Net not only better resemble to the original rainy images in their underlying rain patterns, i.e. the physical factors of rain such as shape and photometry are finely captured by our model, but also achieve more diversity in rain factors, such as orientation, length, width and sparsity. Actually, the diversity in rain factors of TRG-Net is easy to understand, and we can observe from Fig. 7 that TRG-Net has extracted a specific distribution for rain factors, which certainly leads to the diversity in these factors. These results clearly substantiate the superiority of the proposed TRG-Net over the previous method.
5.2.2 The Effectiveness of the Unpaired Generated Data for Deraining
Experiment settings. We adopt PReNet [32], a simple deraining baseline, as the derainer to demonstrate the effectiveness of the generated pseudo train data sets for deraining. The training pairs for PReNet consist of rainless images from the train dataset of the unpaired rain generation experiment and their corresponding pseudo-paired rainy images generated by the rain generators. The competing unpaired methods include the model-based methods, DSC [49] and JCAS [50], and the DL-based methods, CycleGAN [16], DerainCycleGAN [21]and DCD-GAN [22]. Since the DL-based unpaired methods here are all CycleGAN-based framework which can also generate pseudo rainy images, we also train PReNet using the pseudo-paired data generated by these methods, for a fair comparison between these methods and rain generation methods. For convenience, we use the notation $*{}^{*}$ to denote the deraining results of PReNet trained on the pseudo paired data generated by method A. Peak-signal-to-noise (PSNR) and structure similarity (SSIM) are used to quantify the deraining performance, which are calculated in Y channel of YCbCr space following previous works [22, 18].
Experimental results. Table 2 lists the quantitative results of all competing methods on four datasets. We can observe that $*{}^{*}$ outperforms all the other competing methods, which demonstrates the superiority of the unpaired generated rainy images by our model for deraining. Besides, to evaluate the effectiveness of unpaired deraining results for downstream tasks, we perform semantic segmentation experiments using ERFNet [51] on Cityscapes. The mean intersection over union (mIoU) and mean accuracy (mAcc) of the segmentation results are reported in Table 3. Our model Obtains the best segmentation performance among all competing methods, which is significantly superior to the segmentation results of rainy images and shows only a small discrepancy when compared to clean images. In Fig. 9, we provide the deraining results of all competing methods on a representative sample of Cityscapes and their corresponding semantic segmentation results. It can be seen that $*{}^{*}$ consistently achieves evidently superior visual results.
These unpaired experiments demonstrate that the rainy images generated by the proposed rain generation method are not only of higher quality than those generated by the common CCN-based method, but also more superior for deraining and downstream tasks.
5.3 Rain Data Augmentation for Paired Data Set
In this section, we perform rain data augmentation, including in-distribution and out-of-distribution augmentation, for deraining on paired datasets. Specifically, we first conduct in-distribution rain augmentation for deraining on synthetic datasets (Sec. 5.3.1) and real SPA-Data (Sec. 5.3.2) to demonstrate the diversity of generated rain data. Then, OOD rain augmentation, including the orientation and sparsity of rain, are implemented in Sec. 5.3.3 to further validate the controllability superiority of TRG-Net.
5.3.1 Rain Data Augmentation on Synthetic Datasets
Training details of the generator. For a fair comparison with the current SOTA rain generator, VRG-Net [18], we exploit the same discriminator as VRG-Net, i.e., a self-attention discriminator [53] with gradient penalty. The batch size and patch size are set as 10 and $128\times 128$, respectively. The rotatable TV regularizer is adopted for the training of Rain100L and Rain100L-S. The rain generator is trained for 200 epochs on Rain100L and Rain100L-S, and 400 epochs on Rain100H and Rain100H-S. The other training settings are those used in Sec. 5.1. The augmentation rate is set as 0.5 for Rain100L, Rain100L-S and Rain100H-S, and 1% for Rain100H. Note that we input paired clean background $\mathcal{B}$ and rainy image $\mathcal{X}$ for training the rainy image generators, which will lead to better performance than training the model with unpaired samples.
Figure 10: Comparison of deraining results on a test simple from Rain100L-S. (a): The rainy image and groundtruth on Rain100L-S test set. (b)-(e): From top to bottom: the deraining results trained on the original Rain100L-S, augmented data generated by VRG-Net, and augmented data generated by the proposed TRG-Net, respectively, where the deraining methods from left to right are PReNet, SPANet, JORDER_E, RCDNet and DRSformer, respectively.
@@ 5.3.1 Rain Data Augmentation on Synthetic Datasets
Table 5: The generalization performance on SPA-Data. Baseline means the derainers trained on the original dataset without augmentation. VRG-Net and TRG-Net denote augmented training using VRG-Net and the proposed TRG-Net, respectively. The best result is highlighted with bold.
@@
Evaluation on synthetic data. Table 4 provides the deraining results of all competing methods without and with data augmentation, on four datasets. “Baseline” denotes the performance of the derainers trained on the original data set without data augmentation here. As shown in Table 4, the deraining performance of every deep derainer gains a significant improvement with data augmentation in most cases. The performance achieves more improvement on Rain100L-S and Rain100H-S, which accords with human intuition that data augmentation will have more effects when there are less training samples. These results imply that the rainy images generated by the proposed TRG-Net should be closer to those of the original data, which leads to better performance compared to VRG-Net. Fig. 10 visualizes the deraining results of all competing methods on a typical image of Rain100L-S without and with data augmentation. We can observe that the results with TRG-Net augmentation not only more evidently remove rains, but also better preserve the texture and edges of the image.
Table 6: The average PSNR and SSIM of PReNet on the SPA-Data test set. The training set of Baseline is all from original SPA-Data. The training data of VRG-Net and the proposed TRG-Net consist of real pairs randomly selected from SPA-Data and fake pairs generated by different generators. We report the mean of five repeated experiments. The best result is highlighted with bold.
@@
Generalization evaluation on real data. Table 5 shows the generalization performance of models trained on four synthetic datasets and tested on real rainy images of SPA-Data. In this case, the PSNR and SSIM with data augmentation by VRG-Net gain an unsubstantial advantage over the baseline in most cases. By comparison, the performance with data augmentation of TRG-Net still achieves a consistently obvious improvement over the baseline in most test cases. These results demonstrate that the samples generated by our method can improve the deraining performance not only in in-distribution but also in out-of-distribution tasks.
5.3.2 Rain Data Augmentation on Real SPA-Data
@@
Experiment settings. The generator is trained for 400 epochs. The other training settings of TRG-Net are the same as those settings in Sec. 5.3.1. The rain generators are first trained on SPA-Data in a paired manner. Then we take PReNet [32] as the derainer. The training data for the derainer consist of a small number of pairs (i.e., 1K and 2K) randomly selected from SPA-Data and $N$K fake pairs generated by the generators. For comparison, the same number of paired data from SPA-Data is also randomly chosen as the training set of the derainer, which is called “Baseline” in this experiment.
The experimental results. The PSNR and SSIM on the test set of SPA-Data are shown in Table 6888To avoid randomness of the experiment, we report the mean of five repetitive experiments following [18].. From Table 6, we can observe that better performance can be achieved with generated data samples than that with even more original data samples [^9] This can be rationally explained by the fact that the generated rainy images are with much less repetitively but possibly with a more diversity (possibly even generating new rain patterns that not appear in the original training data, as the previous analysis) in a highly controllable manner than the original training data.. In addition, the derainer trained on the samples generated by the proposed TRG-Net obtains a better result than those trained on the same and even more number of data samples generated by the VRG-Net [18]. These results validate the superior quality of the generated samples by the proposed TRG-Net beyond those generated by VRG-Net.
Table 7: The PSNR and SSIM comparison on test data with orientation degree of $[-30^{\circ},30^{\circ}]$ and ±$\pm[30^{\circ},60^{\circ}]$. Baseline represents the deraining results of PReNet [32] trained on original training set. VRG-Net, TRG-Net${}_{[-30^{\circ},30^{\circ}]}$ and TRG-Net${}_{[-60^{\circ},60^{\circ}]}$ are the deraining results of PReNet augmented on VRG-Net, TRG-Net${}_{[-30^{\circ}\!,30^{\circ}]}$ and TRG-Net${}_{[-60^{\circ}\!,60^{\circ}]}$, respectively.
@@
Figure 11: The visual deraining results of a typical image in the test rainy images with orientation degrees ±$\pm[30^{\circ},60^{\circ}]$. Upper row: from left to right are the input rainy image, deraining results of DSC [49], JCAS [50] and PReNet [32] without data augmentation, respectively. Lower row: from left to right are the groundtruth, deraining results of PReNet augmented on VRG-Net [18], TRG-Net${}_{[-30^{\circ},30^{\circ}]}$, TRG-Net${}_{[-60^{\circ},60^{\circ}]}$, respectively.
Table 8: The deraining performance on Cityscapes with varying levels of sparsity. Baseline represents the deraining results of PReNet trained on original training set. VRG-Net, TRG-Net are the deraining results of PReNet through in-distribution augmentation on VRG-Net and TRG-Net, respectively. ${}^{\dagger}$ denotes the deraining results of augmented PReNet on out-of-distribution rain by controlling the sparsity using TRG-Net. The best result is highlighted with bold.
out-of-distribution in-distribution out-of-distribution
5mm/hr 25mm/hr 50mm/hr 75mm/hr 100mm/hr
Methods PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
Baseline 34.20 0.995 28.70 0.990 23.58 0.976 21.42 0.963 20.21 0.952
VRG-Net[18] 33.24 0.995 28.20 0.990 23.58 0.976 21.35 0.962 20.10 0.950
TRG-Net 34.33 0.995 29.00 0.990 24.75 0.979 21.88 0.964 20.48 0.952
${}^{\dagger}$ 34.53 0.995 29.15 0.990 24.71 0.979 22.00 0.966 20.58 0.954
5.3.3 OOD Rain Augmentation
To further verify the superiority of our controllable rain generation, we generate OOD rain for OOD deraining by controlling the rain factors of trained TRG-Net. Specifically, the two main factors influencing the appearance of rain, namely orientation and sparsity, are augmented in the following.

### v2

@@ V-A Rain Generator Verification
The learned distributions [^7] The distribution is approximated by the histogram obtained by sampling 10000 times. of rain factors in Rain100L are shown in Fig. 7, which can be easily used to further generate different rain types by adjusting rain factors in a controllable manner. As shown in Fig. 7 (a), the learned orientation degree $\theta$ distribution is close to a Gaussian distribution, with values mostly between $-40$ and $40$, which complies with the intuitive orientation distribution in Rain100L [5]. By model (21), we can manually input the orientation degree into the trained generator to control the orientation of the generated rain steaks. Fig. 7 (a) shows the rain generated with input $\theta$ of $0^{\circ}$, $30^{\circ}$, $60^{\circ}$, $-30^{\circ}$, respectively. It should be noted that this manner offers the capability to generate OOD rains, e.g., the $60^{\circ}$ rain in Fig. 7 (a), which will be beneficial for deraining on OOD rain data sets (see OOD rain augmentation in Sec. V-C3). Yet this is difficult to be achieved by traditional DL-based generators, since black-box model can hardly generalize samples beyond the distribution of training data. Similarly, in Fig. 7 (b), (c) and (d), we can observe that the proposed method can also learn the distributions of $s_{l}$, $s_{w}$ and $\tau$, respectively. We can control rain from thin to thick and from short to long by scaling the rain kernels according to different $s_{w}$ and $s_{l}$ values, respectively, and control the sparsity of rain by adjusting the input of $\tau$.
V-B Unpaired Rain Generation
@@
V-B1 Unpaired Rain Generation
TABLE I: The FID and KID of unpaired rain generation on four rain datasets. Bold indicates the best result.
@@ V-B1 Unpaired Rain Generation
Figure 8: The results of unpaired rain generation by VRG-Net and the proposed TRG-Net on three synthetic datasets and a real dataset. (a)-(d) are the background images, the rainy images generated by VRG-Net, the rainy images generated by the proposed TRG-Net and the reference rainy images which are used for unpaired training, respectively. From top to bottom are Rain100L, Cityscapes, Kitti and SPA-Data datasets, respectively. Every dataset displays two generated samples.
Experiment settings. We perform unpaired rain generation on three synthetic rain datasets, Rain100L, Cityscapes [44] and Kitti888 The rains in Cityscapes and Kitti are synthesised [44] with varying rainfall rate to evaluate the effect of rain on outdoor computer vision tasks. For them, we select the 100mm/hr rain as training set and test set in this experiment. As the rain orientations in Cityscapes and Kitti are correlated with the position, we have further included three learnable masks on the rain maps to distinguish different orientations in different positions. [44], and a real rain dataset, SPA-Data [26], to verify the superiority of our rain generator. Specifically, to construct a complete unpaired training dataset, we remove the rainless images of half of the data samples and remove the rainy images of the other half of the data samples. The rotatable TV regularizer is used for the training of Rain100L. The generator is trained for 400 epochs on Cityscapes and Kitti, and 30 epochs and 250 epochs on Rain100L and SPA-Data, respectively. The same patch-based discriminator [17] is employed for all generators. Inspired by [47], we use the high frequency components of the rainy images instead of the rainy image itself as input to the discriminator for better discrimination [^9] Please refer to the supplementary material for more details. The other training settings are the same as those used in Sec. V-A. We utilize Fréchet Inception Distance (FID) and Kernel Inception Distance (KID) as metrics for assessing the quality of generated rainy images, where lower values indicate better performance.
TABLE II: The quantitative results of all competing methods on synthetic and real datasets. A indicates the deraining results of PReNet [30] trained on the pseudo-paired data generated by method A. The best result is highlighted with bold.
@@
CycleGAN [16] 25.11 0.832 22.62 0.908 26.88 0.937 22.71 0.866
DerainCycleGAN [22] 28.50 0.910 21.10 0.871 22.31 0.938 22.01 0.863
DCD-GAN [23] 22.48 0.821 19.84 0.850 23.55 0.903 19.81 0.786
VRG-Net [18] 27.84 0.855 21.94 0.768 21.08 0.906 34.81 0.931
TRG-Net 32.33 0.949 24.86 0.959 32.15 0.983 35.80 0.945
@@
Experimental results. Table I shows the FID and KID results on four datasets. Our model achieves lower FID and KID compared to VRG-Net on all datasets, expecting the KID on Kitti, which means that the rainy images generated by the proposed TRG-Net has a higher rain quality than VRG-Net [18] and is able to capture the rain generation mechanism under different environment conditions. In Fig. 8, we show some typical results of the unpaired rain generation. It is easy to observe that the rain streaks generated by VRG-Net, which represents the CNN-based method, could be largely lacking in diversity, for example, the rain streaks in two random generation results on different background images are very similar, as shown in Fig. 8 (b). In contrast, the rainy images generated by the proposed TRG-Net not only better resemble to the original rainy images in their underlying rain patterns, i.e. the physical factors of rain such as shape and photometry are finely captured by our model, but also achieve more diversity in rain factors, such as orientation, length, width and sparsity. Actually, the diversity in rain factors of TRG-Net is easy to understand, and we can observe from Fig. 7 that TRG-Net has extracted a specific distribution for rain factors, which certainly leads to the diversity in these factors. These results clearly substantiate the superiority of the proposed TRG-Net over the previous method.
V-B2 The Effectiveness of the Unpaired Generated Data for Deraining
Experiment settings. We adopt PReNet [30], a simple deraining baseline, as the derainer to demonstrate the effectiveness of the generated pseudo train data sets for deraining. The training pairs for PReNet consist of rainless images from the train dataset of the unpaired rain generation experiment and their corresponding pseudo-paired rainy images generated by the rain generators. The competing unpaired methods include the model-based methods, DSC [48] and JCAS [49], and the DL-based methods, CycleGAN [16], DerainCycleGAN [22]and DCD-GAN [23]. Since the DL-based unpaired methods here are all CycleGAN-based framework which can also generate pseudo rainy images, we also train PReNet using the pseudo-paired data generated by these methods, for a fair comparison between these methods and rain generation methods. For convenience, we use the notation A to denote the deraining results of PReNet trained on the pseudo paired data generated by method A. Peak-signal-to-noise (PSNR) and structure similarity (SSIM) are used to quantify the deraining performance, which are calculated in Y channel of YCbCr space following previous works [23, 18].
Experimental results. Table II lists the quantitative results of all competing methods on four datasets. We can observe that TRG-Net outperforms all the other competing methods, which demonstrates the superiority of the unpaired generated rainy images by our model for deraining. Besides, to evaluate the effectiveness of unpaired deraining results for downstream tasks, we perform semantic segmentation experiments using ERFNet [50] on Cityscapes. In Fig. 9, we provide the deraining results of all competing methods on a representative sample of Cityscapes and their corresponding semantic segmentation results [^10] The quantitative results of semantic segmentation are presented in the supplementary material due to space limitations.. It can be seen that TRG-Net not only achieves clearly superior visual results in deraining, but also in segmentation.
TABLE III: The deraining results on synthetic datasets. Baseline means the derainers trained on the original dataset without augmentation. VRG-Net and TRG-Net denote augmented training using VRG-Net and the proposed TRG-Net, respectively. Bold is the best result.
@@ V-B2 The Effectiveness of the Unpaired Generated Data for Deraining
TABLE IV: The generalization performance on SPA-Data. Baseline means the derainers trained on the original dataset without augmentation. VRG-Net and TRG-Net denote augmented training using VRG-Net and the proposed TRG-Net, respectively. The best result is highlighted with bold.
@@
V-C Rain Data Augmentation for Paired Data Set
In this section, we perform rain data augmentation, including in-distribution and out-of-distribution augmentation, for deraining on paired datasets. Specifically, we first conduct in-distribution rain augmentation for deraining on synthetic datasets (Sec. V-C1) and real SPA-Data (Sec. V-C2) to demonstrate the diversity of rain generated by the proposed method. Then, OOD rain augmentation are implemented in Sec. V-C3 to further validate the controllability superiority of TRG-Net.
V-C1 Rain Data Augmentation on Synthetic Datasets
Training details of the generator. For a fair comparison with VRG-Net [18], we exploit the same discriminator as VRG-Net, i.e., a self-attention discriminator [51] with gradient penalty. The batch size and patch size are set as 10 and $128\times 128$, respectively. The rotatable TV regularizer is adopted for the training of Rain100L and Rain100L-S. The rain generator is trained for 200 epochs on Rain100L and Rain100L-S, and 400 epochs on Rain100H and Rain100H-S. The other training settings are those used in Sec. V-A. The augmentation rate is set as 0.5 for Rain100L, Rain100L-S and Rain100H-S, and 1% for Rain100H.
Evaluation on same-domain datasets. Table III provides the deraining results of all competing methods without and with data augmentation, on four datasets. “Baseline” denotes the performance of the derainers trained on the original data set without data augmentation here. As shown in Table III, the deraining performance of every deep derainer gains a significant improvement with data augmentation in most cases. The performance achieves more improvement on Rain100L-S and Rain100H-S, which accords with human intuition that data augmentation will have more effects when there are less training samples. These results imply that the rainy images generated by the proposed TRG-Net should be closer to those of the original data, which leads to better performance compared to VRG-Net.
Evaluation on cross-domain datasets. To further validate the diversity of the samples generated by rain generators in the paired manner, we test the generalization performance of models trained on four augmented datasets. The performance on real rainy images of SPA-Data is shown in Table IV. In this case, the PSNR and SSIM with data augmentation by VRG-Net gain an unsubstantial advantage over the baseline in most cases. By comparison, the performance with data augmentation of TRG-Net still achieves a consistently obvious improvement over the baseline in most test cases. These results demonstrate that the samples generated by our method are diverse, which can improve the deraining performance not only in in-distribution but also in out-of-distribution tasks.
TABLE V: The average PSNR and SSIM of PReNet on the SPA-Data test set. The training set of Baseline is all from original SPA-Data. The training data of VRG-Net and the proposed TRG-Net consist of real pairs randomly selected from SPA-Data and fake pairs generated by different generators. We report the mean of five repeated experiments. The best result is highlighted with bold.
@@ V-C1 Rain Data Augmentation on Synthetic Datasets
V-C2 Rain Data Augmentation on Real SPA-Data
@@
Experiment settings. The generator is trained for 400 epochs. The other training settings of TRG-Net are the same as those settings in Sec. V-C1 of the main text. The rain generators are first trained on SPA-Data in a paired manner. Then we take PReNet [30] as the derainer. The training data for the derainer consist of a small number of pairs (i.e., 1K and 2K) randomly selected from SPA-Data and $N$K fake pairs generated by the generators. For comparison, the same number of paired data from SPA-Data is also randomly chosen as the training set of the derainer, which is called “Baseline” in this experiment.
The experimental results. The PSNR and SSIM on the test set of SPA-Data are shown in Table V [^11] To avoid randomness of the experiment, we report the mean of five repetitive experiments following [18].. From Table V, we can observe that better performance can be achieved with generated data samples than that with even more original data samples In addition, the derainer trained on the samples generated by the proposed TRG-Net obtains a better result than those trained on the same and even more number of data samples generated by the VRG-Net [18]. These results validate the superior quality of the generated samples by the proposed TRG-Net beyond those generated by VRG-Net.
TABLE VI: The PSNR and SSIM comparison on test data with orientation degree of $[-30^{\circ},30^{\circ}]$ and ±$\pm[30^{\circ},60^{\circ}]$. Baseline represents the deraining results of PReNet [30] trained on original training set. VRG-Net, TRG-Net${}_{[-30^{\circ},30^{\circ}]}$ and TRG-Net${}_{[-60^{\circ},60^{\circ}]}$ are the deraining results of PReNet augmented on VRG-Net, TRG-Net${}_{[-30^{\circ}\!,30^{\circ}]}$ and TRG-Net${}_{[-60^{\circ}\!,60^{\circ}]}$, respectively.
@@
V-C3 OOD Rain Augmentation
To further verify the superiority of our controllable rain generation, we generate OOD rain for OOD deraining by controlling the rain factors of trained TRG-Net. Specifically, the rain orientation augmentation are conducted in the following [^12] We also perform the rain sparsity augmentation experiment in supplementary material..